OpenStudy (anonymous):

8a^2b^3 -------- 3a^2b^3 how would you do this?

OpenStudy (anonymous):

cancel everything, get \[\frac{8}{3}\]

OpenStudy (cwrw238):

a^2 / a^2 = 1 b^3 / b^3 = 1 so its 8/3

OpenStudy (anonymous):

since \[\frac{a^2}{a^2}=1\] and likewise \[\frac{b^3}{b^3}=1\]

OpenStudy (anonymous):

thanks!